OpenStudy (anonymous):

How do I find the units digit of 2012^103

OpenStudy (asnaseer):

HINT: \(2012=2^2\times503\). Look for patterns in the last digit of \(2^n\) and \(503^n\) as \(n=1,2,...\) then use this information to deduce the result.

OpenStudy (anonymous):

still confused can you do a step by step explaination by any chance?

OpenStudy (tkhunny):

2 (2) 2^2 = 4 (4) 2^3 = 8 (8) 2^4 = 16 (6) 2^5 = 32 (2) 2^6 = 64 (4) Do you believe that this pattern might continue? 2, 4, 8, 6, 2, 4, 8, 6, etc.

OpenStudy (anonymous):

Yes sorry i was away from my computer

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

Can someone please help!

OpenStudy (tkhunny):

Did help. So did asnaseer. What are your thoughts concerning these hints?

OpenStudy (anonymous):

Yes a little. So I notice the pattern is {2,4,8,16} So I take the exponent and divide by 4. I get 25.75. Then I multiply it by the divisor(4) and I get 100. Then I subtract 100 from 103 to get 3. 3 is the amount of moves I make around the mod circle which lands on 6. So the units digit is 6?

OpenStudy (anonymous):

 IS this correct?